# Emergence of rogue-like waves in a reaction-diffusion system: Stochastic output from deterministic dynamics

### v1

@@ Acknowledgements.
We are grateful to A. van Kan for helpful comments. This work was supported in parts by the National Science Foundation under grant DMS-1908891 (EK), by the Israel Science Foundation under grant 1224/21 (AY), and by grant 2022072 from the United States - Israel Binational Science Foundation (BSF), Jerusalem, Israel (EK and AY).
@@ Appendix A Fitting of distributions
Figure 6: Quantiles of $\Delta t$ (a) and $\Delta x$ (b) for the exponential and Gamma ($\Gamma$) distributions (in $\bullet$), respectively, while the insets show the PDFs [from Fig. 5(c)] on logarithmic scale. Dashed lines represent perfect distributions while solid lines represent a linear fit with the coefficient of determination $R^{2}$.

### v2

@@ Acknowledgements.
We are grateful to A. van Kan and N. Smith for helpful comments. This work was supported in parts by the National Science Foundation under grant DMS-1908891 (EK), by the Israel Science Foundation under grant 1224/21 (AY), and by grant 2022072 from the United States - Israel Binational Science Foundation (BSF), Jerusalem, Israel (EK and AY).
@@ Appendix A Fitting of distributions
Figure 6: Quantiles of (a) $\Delta t$ and (b) $\Delta x$ for the exponential and Gamma ($\Gamma$) distributions (in $\bullet$), respectively, while the insets show the PDFs [from Fig. 5(b)] on a logarithmic scale. Dashed lines represent perfect distributions while solid lines represent a linear fit with regression coefficient $R^{2}$.